# Retrieval-Augmented Generation for Knowledge-Intensive NLP Tasks

### 22 May 2020 v1

@@ 5 Related Work
Our document index can be seen as a large external memory for neural networks to attend to, analagous to memory networks [58, 48]. Concurrent work [14] learns to retrieve a trained embedding for each entity in the input, rather than to retrieve raw text as in our work. Other work improves the ability of dialog models to generate factual text by attending over fact embeddings [9, 13] or, closer to our work, over retrieved text directly [15]. A key feature of our memory is that it is comprised of raw text rather distributed representations, which makes the memory both (i) human-readable, lending a form of interpretability to our model, and (ii) human-writable, enabling us to dynamically update the model’s memory by editing the document index.
@@ 6 Discussion
In this work, we presented hybrid generation models with access to parametric and non-parametric retrieval-based external memory, in the form of Wikipedia. We showed that our RAG models obtain state-of-the-art performance on open domain question answering. We found that people prefer RAG’s generation over purely parametric BART and find RAG more factual, and we conducted a detailed investigation of the learned retrieval component, validating its effectiveness. We also showed that the model’s grounding in external data leads it to generate more diverse, and illustrated by how the retrieval index can be hot-swapped on the fly without having to retrain the model. In future work, it would be interesting to investigate if the two components can be jointly pre-trained from scratch, either on a denoising objective similar to BART, or through some other objective. Our work opens new research directions on how parametric and non-parametric memories interact and how to most effectively combine the different components, showing promise in being applied to a wide variety of NLP tasks.
@@ Acknowledgments
EP thanks supports from the NSF Graduate Research Fellowship.
Appendix A Human evaluation
@@
Appendix B Further details on Open-Domain QA
@@
Appendix C Further details on FEVER
@@
Appendix D "Null document" Probabilities
@@
Appendix E Parameters
Our RAG models contain the trainable parameters for the BERT-base query and document encoder of DPR, with 110M parameters each (although we do not train the document encoder ourselves) and 406M trainable parameters from BART-large, 406M parameters, making a total of 626M trainable parameters. The best performing "closed-book" (parametric only) open-domain QA model is T5-11B with 11 Billion trainable parameters. The T5 model with the closest number of parameters to our models is T5-large (770M parameters), which achieves a score of 28.9 EM on Natural Questions [46], substantially below the 44.5 that RAG-Sequence achieves, indicating that hybrid parametric/non-parametric models require far fewer trainable parameters for strong open-domain QA performance. The non-parametric memory index does not consist of trainable parameters, but does consists of 21M 728 dimensional vectors, consisting of 15.3B values.
Appendix F Retrieval Collapse

### 07 Dec 2020 v2

@@ 5 Related Work
Our document index can be seen as a large external memory for neural networks to attend to, analogous to memory networks [64, 55]. Concurrent work [14] learns to retrieve a trained embedding for each entity in the input, rather than to retrieve raw text as in our work. Other work improves the ability of dialog models to generate factual text by attending over fact embeddings [9, 13] or, closer to our work, over retrieved text directly [15]. A key feature of our memory is that it is comprised of raw text rather distributed representations, which makes the memory both (i) human-readable, lending a form of interpretability to our model, and (ii) human-writable, enabling us to dynamically update the model’s memory by editing the document index.
Retrieve-and-Edit approaches
Our method shares some similarities with retrieve-and-edit style approaches, where a similar training input-output pair is retrieved for a given input, and then edited to provide a final output. These approaches have proved successful in a number of domains including Machine Translation  [18, 22] and Semantic Parsing [21]. Our approach does have several differences, including less of emphasis on lightly editing a retrieved item, but on aggregating content from several pieces of retrieved content, as well as learning latent retrieval, and retrieving evidence documents rather than related training pairs. This said, RAG techniques may work well in these settings, and could represent promising future work.
@@ 6 Discussion
In this work, we presented hybrid generation models with access to parametric and non-parametric memory. We showed that our RAG models obtain state of the art results on open-domain QA. We found that people prefer RAG’s generation over purely parametric BART, finding RAG more factual and specific. We conducted an thorough investigation of the learned retrieval component, validating its effectiveness, and we illustrated how the retrieval index can be hot-swapped to update the model without requiring any retraining. In future work, it may be fruitful to investigate if the two components can be jointly pre-trained from scratch, either with a denoising objective similar to BART or some another objective. Our work opens up new research directions on how parametric and non-parametric memories interact and how to most effectively combine them, showing promise in being applied to a wide variety of NLP tasks.
Broader Impact
This work offers several positive societal benefits over previous work: the fact that it is more strongly grounded in real factual knowledge (in this case Wikipedia) makes it “hallucinate” less with generations that are more factual, and offers more control and interpretability. RAG could be employed in a wide variety of scenarios with direct benefit to society, for example by endowing it with a medical index and asking it open-domain questions on that topic, or by helping people be more effective at their jobs.
With these advantages also come potential downsides: Wikipedia, or any potential external knowledge source, will probably never be entirely factual and completely devoid of bias. Since RAG can be employed as a language model, similar concerns as for GPT-2 [50] are valid here, although arguably to a lesser extent, including that it might be used to generate abuse, faked or misleading content in the news or on social media; to impersonate others; or to automate the production of spam/phishing content [54]. Advanced language models may also lead to the automation of various jobs in the coming decades [16]. In order to mitigate these risks, AI systems could be employed to fight against misleading content and automated spam/phishing.
@@ Acknowledgments
The authors would like to thank the reviewers for their thoughtful and constructive feedback on this paper, as well as HuggingFace for their help in open-sourcing code to run RAG models. The authors would also like to thank Kyunghyun Cho and Sewon Min for productive discussions and advice. EP thanks supports from the NSF Graduate Research Fellowship. PL is supported by the FAIR PhD program.
Appendices for Retrieval-Augmented Generation for Knowledge-Intensive NLP Tasks
Appendix A Implementation Details
Appendix B Human Evaluation
@@
Appendix C Training setup Details
We train all RAG models and BART baselines using Fairseq [45]. [^2] https://github.com/pytorch/fairseq We train with mixed precision floating point arithmetic [40], distributing training across 8, 32GB NVIDIA V100 GPUs, though training and inference can be run on one GPU. We find that doing Maximum Inner Product Search with FAISS is sufficiently fast on CPU, so we store document index vectors on CPU, requiring $\sim 100$ GB of CPU memory for all of Wikipedia. After submission, We have ported our code to HuggingFace Transformers [66] [^3] https://github.com/huggingface/transformers, which achieves equivalent performance to the previous version but is a cleaner and easier to use implementation. This version is also open-sourced. We also compress the document index using FAISS’s compression tools, reducing the CPU memory requirement to 36GB. Scripts to run experiments with RAG can be found at https://github.com/huggingface/transformers/blob/master/examples/rag/README.md and an interactive demo of a RAG model can be found at https://huggingface.co/rag/
Appendix D Further Details on Open-Domain QA
@@
CuratedTrec preprocessing
The answers for CuratedTrec are given in the form of regular expressions, which has been suggested as a reason why it is unsuitable for answer-generation models [20]. To overcome this, we use a pre-processing step where we first retrieve the top 1000 documents for each query, and use the answer that most frequently matches the regex pattern as the supervision target. If no matches are found, we resort to a simple heuristic: generate all possible permutations for each regex, replacing non-deterministic symbols in the regex nested tree structure with a whitespace.
@@
Appendix E Further Details on FEVER
@@
Appendix F Null Document Probabilities
@@
Appendix G Parameters
Our RAG models contain the trainable parameters for the BERT-base query and document encoder of DPR, with 110M parameters each (although we do not train the document encoder ourselves) and 406M trainable parameters from BART-large, 406M parameters, making a total of 626M trainable parameters. The best performing "closed-book" (parametric only) open-domain QA model is T5-11B with 11 Billion trainable parameters. The T5 model with the closest number of parameters to our models is T5-large (770M parameters), which achieves a score of 28.9 EM on Natural Questions [52], substantially below the 44.5 that RAG-Sequence achieves, indicating that hybrid parametric/non-parametric models require far fewer trainable parameters for strong open-domain QA performance. The non-parametric memory index does not consist of trainable parameters, but does consists of 21M 728 dimensional vectors, consisting of 15.3B values. These can be easily be stored at 8-bit floating point precision to manage memory and disk footprints.
Appendix H Retrieval Collapse
@@
Appendix I Number of instances per dataset
The number of training, development and test datapoints in each of our datasets is shown in Table 7.
Table 7: Number of instances in the datasets used. *A hidden subset of this data is used for evaluation
Task Train Development Test
Natural Questions 79169 8758 3611
TriviaQA 78786 8838 11314
WebQuestions 3418 362 2033
CuratedTrec 635 134 635
Jeopardy Question Generation 97392 13714 26849
MS-MARCO 153726 12468 101093*
FEVER-3-way 145450 10000 10000
FEVER-2-way 96966 6666 6666